STATISTICS. please help.
2.An automatic soda dispenser is regulated so that it dispenses an average of 200 ml per cup. Suppose the amount of drink dispensed is normally distributed with a standard deviation equal to 15 ml.
a.What is the probability that a cup contains between 191 and 209 ml?
b.What is the probability that a cup will overflow if 230 ml cups are used?

Answer by Boreal(15235)

z=(x-mean(/sd
z=(191-200)/15=-0.6
z=(209-200)/15=+0.6
the probability of z between those values is 0.4515
----
z=(230-200)/15=+2
probability z > 2 is 0.0228